What is the slope of the line with equation 4x+6y=24?
Answer by fractalier(6550)

If you solve for y, you'll automatically put the equation in slope-intercept form (y = mx + b) from which you can read the slope...let me show you...
From 4x + 6y = 24, subtract 4x from both sides
6y = -4x + 24
Then divide both sides by six and reduce where possible
y = -4/6 x + 24/6
y = -2/3 x + 4
Hence the slope is -2/3.